OpenStudy (anonymous):

Find the equation of a straight line through points (2,6) parallel to the line with equation 2x+5y=1

OpenStudy (mathlegend):

Okay first you want to get that equation into slope intercept form y=mx+b

OpenStudy (mathlegend):

Can you do that?

OpenStudy (mathlegend):

2x+5y=1 How do you get the 2x to the other side?

OpenStudy (anonymous):

so 5y = 1-2x?

OpenStudy (mathlegend):

Well remember we want it in a specific order... y=mx+b 5y=-2x+1

OpenStudy (mathlegend):

Now how we get rid of the '5y' which is being multiplied?

OpenStudy (anonymous):

y= (-2x+1)/5?

OpenStudy (mathlegend):

Okay so we have... \[\frac{ -2 }{ 5}x+\frac{ 1 }{ 5}\]

OpenStudy (mathlegend):

So now we need to use the distance formula for the Coordinate points. y-y1=m(x-x1)

OpenStudy (mathlegend):

Can you plug in the numbers?

OpenStudy (anonymous):

y-6=-2(x-2)?

OpenStudy (mathlegend):

Remember our slope was \[\frac{ -2 }{ 5}\]

OpenStudy (mathlegend):

y-6= -2/5(x-2)

OpenStudy (mathlegend):

Now when you distribute what do you get?

OpenStudy (anonymous):

y=-2/5x\[-6\frac{ 4}{ 5 }\]?

OpenStudy (mathlegend):

y-6=-2/5x+4/5 That is what you should have gotten.

OpenStudy (anonymous):

ohh, sorry i put the -6 to the other side, but forgot to make it positive instead of negative

OpenStudy (mathlegend):

So now lets get the y by itself. So lets add 6 to both sides. Tell me what you get once you add 6 to both sides.

OpenStudy (anonymous):

y=-2/5x+34/5?

OpenStudy (mathlegend):

Good that is your answer.

OpenStudy (anonymous):

You truly are a math legend :) thankyou :)

OpenStudy (mathlegend):

No problem :)